# Pion Valence Structure at Intermediate x in the Residual Field Model

### v1

@@ III.2 Hadronic Tensor and The Scattering Amplitude
$(2\pi)^{4}\delta^{(4)}(p_{\pi}+q-\sum_{i=1}^{3}k_{i}-k_{R})A^{\mu\dagger}A^{\nu}$ (10)
where the sum $\sum_{\{h_{i},\tau_{i}\}}$ is over all the spins of the valence quarks and the residual system. The Lorentz invariant phase space measure in light front coordinates is,
@@
$=$ $\frac{dk^{+}d^{2}\mathbf{k}_{\perp}}{2k^{+}}|_{k^{-}=\frac{k_{\perp}^{2}+m^{2}}{k^{+}}}=\frac{dxd^{2}\mathbf{k}_{\perp}}{2x}|_{k^{-}=\frac{k_{\perp}^{2}+m^{2}}{xP_{\pi}^{+}}}$
@@
×$\times\delta^{(2)}(\mathbf{p_{\pi}^{\perp}}+\mathbf{q}^{\perp}-\mathbf{k^{\prime\perp}_{1}}-\mathbf{k_{2}^{\perp}}-\mathbf{k}_{R}^{\perp}).$
Taking into account of Eq.(7) according to which $q^{+}=0$ and $k^{\prime+}_{1}=k_{1}^{+}$ one obtains:
@@
where $x_{1}={k_{1}^{+}\over p_{\pi}^{+}}$, $x_{2}={k_{2}^{+}\over p_{\pi}^{+}}$, and $x_{R}={k_{R}^{+}\over p_{\pi}^{+}}$. Using $k_{1}^{\prime}=p_{\pi}-k_{2}-k_{3}-k_{R}$ and neglecting initial transverse momenta compared to $q_{\perp}$,i.e. $k_{1,\perp}^{\prime 2}=(\mathbf{k}_{1,\perp}+\mathbf{q}_{\perp})^{2}\sim\mathbf{q}_{\perp}^{2}\gg k_{1,\perp}^{2}$ as well as using the condition $p_{\pi}^{+}\gg m_{\pi},m,\ m_{R},\ k_{1,2}^{\perp},\ k_{R}$ one evaluates:
@@
$\approx$ $\delta\left(\frac{2p_{\pi}\cdot q}{p_{\pi}^{+}}-\frac{Q^{2}}{x_{1}p_{\pi}+}\right)=\frac{x_{1}p_{\pi}^{+}}{2p_{\pi}\cdot q}\delta(x_{1}-x_{B}),$
where Bjorken x is defined as $x_{B}=\frac{Q^{2}}{2P\cdot q}$. Finally with the same conditions:
$\delta(p_{\pi}^{+}+q^{+}-k^{\prime+}_{1}-k^{+}_{2}-k_{R}^{+})=\delta(k^{+}_{1}+k^{+}_{2}+k_{R}^{+})$ (15)
@@
$W^{\mu\nu}(x,Q^{2})=\frac{1}{4m_{\pi}}\sum_{\{h_{i},\tau_{i}\}}\int[dx][d^{2}\mathbf{k}_{\perp}]{x_{1}\over p_{\pi}q}\delta(x_{1}-x_{B})A^{\mu\dagger}A^{\nu}$ (16)
@@
$\left[d^{2}\mathbf{k}_{\perp}\right]$ $=$ $16\pi^{3}\delta^{(2)}(\sum_{i=1}^{2}\mathbf{k}_{i,\perp}+\mathbf{k}_{R,\perp})\frac{d^{2}\mathbf{k}_{R,\perp}}{16\pi^{3}}\prod_{i=1}^{2}\frac{d^{2}\mathbf{k}_{i,\perp}}{16\pi^{3}}$ (17)
@@ III.3 Scattering Amplitude in Residual Field Model
To calculate the amplitude presented in Fig.(1(b)) we apply diagrammatic rules of Light-Front perturbation theory. Since momentum transferred to the quark/anti-quark in the pion is significantly larger than their initial momenta, the interference terms will be small and the cross section will represent the incoherent sum of the square of amplitudes corresponding to the scattering from quark and anti-quark. Thus we first calculate amplitude of scattering from quark and then use the result to evaluate the scattering from the anti-quark.
To calculate the amplitude we noticed that the diagram of Fig.(1(b)) is characterized by two intermediate states (indicated by vertical dashed lines) that are defined by two light-front propagators. With the momenta assigned for each particles in the scattering defined in the diagram one obtains:
$A^{\mu}(k_{1}^{\prime},h_{1}^{\prime},k_{2},h_{2},k_{R},h_{R})=\sum\limits_{h_{V}}{\bar{u}(k_{1}^{\prime},h_{1}^{\prime})ie_{q}\gamma^{\mu}(/\hskip-5.0ptk_{1}+m_{q})v(k_{2},h_{2})\Gamma^{V\to q\bar{q}}\chi_{V}\over k_{1}^{+}D_{2}}{\chi_{V}^{\dagger}\chi_{R}^{\dagger}\Gamma^{\pi\to VR}\chi_{\pi}\over k_{V}^{+}D_{1}},$ (18)
where $\Gamma^{\pi\rightarrow VR}$ is the effective vertex for the transition of pion into valence $q\bar{q}$ cluster and $\Gamma^{V\rightarrow q\bar{q}}$ describes the transition f the cluster to valence quark and anti-quark. Also, $h_{i}$ is the helicity of the i’th valence parton and $h_{1}^{\prime}$ is the helicity of scattered quark.
For the denominators $D_{1}$ and $D_{2}$ according to LF diagrammatic rules, one has:
@@
Furthermore, we use LF sum rule for the numerator of the propagator of off-energy shell quark with momentum $k_{1}$:
$(/\hskip-5.0ptk_{1}+m_{q})=\sum\limits_{h1}u(k_{1},h_{1})\bar{u}(k_{1},h_{1})+{\gamma^{+}(k^{2}-m^{2})\over k^{+}}$ (21)
and notice that since we are interested in the $A^{+}$ component of the scattering amplitude, but due to the relation $\gamma^{+}\gamma^{+}=0$, the second part of Eq.(21) does not contribute to the $+$ component of the amplitude. Inserting the first part of Eq.(21) into Eq.(18) and introducing phenomenological LF wave functions for valence $q\bar{q}$ cluster:
@@
and residual system
$\psi_{q\bar{q}}(\beta_{1},\beta_{2},\mathbf{k}_{1,\perp},\mathbf{k}_{2,\perp})=\frac{\bar{u}(k_{1},h_{1})\Gamma^{V\rightarrow q\bar{q}}v(k_{2},h_{2})\chi_{V}}{m^{2}_{V}-\sum_{i=1}^{2}\frac{k^{2}_{i,\perp}+m^{2}_{i}}{\beta_{i}}}$ (23)
For the ”+”-component of the scattering amplitude assuming that the residual state has zero helicity and writing explicitly the total zero helicity state of the valence $q\bar{q}$ state as $\xi_{q\bar{q}}={u(k_{1},h_{1}){\bar{u}}(k_{2},-h_{1})-u(k_{1},-h_{1})\bar{u}(k_{2},h_{1})\over\sqrt{2}}$ one obtains:
$A^{+}(k_{1}^{\prime},h_{1}^{\prime},k_{2},-h_{1},k_{R})=\sum\limits_{h_{1}}\left[\bar{u}(k_{1}^{\prime},h_{1}^{\prime})ie_{q}\gamma^{+}u(k_{1},h_{1}){\psi_{q\bar{q}}(\beta_{1},\beta_{2},\mathbf{k}_{1,\perp},\mathbf{k}_{2,\perp})\over\sqrt{2}\beta_{1}}\right.$
@@
By neglecting the masses of valence quarks, one evaluates only helicity conserving term of $\bar{u}(k_{1}^{\prime},h_{1}^{\prime})\gamma^{+}u(k_{1},h_{1})=\bar{v}(k_{1},h_{1})\gamma^{+}v(k_{1}^{\prime},h_{1}^{\prime})=2k_{1}^{+}\delta^{h^{\prime}_{1},h_{1}}$. This results in:
$A^{+}(k_{1}^{\prime},h_{1},k_{2},-h_{1},k_{R})=\left[2k_{1}^{+}ie_{q}{\psi_{q\bar{q}}(\beta_{1},\beta_{2},\mathbf{k}_{1,\perp},\mathbf{k}_{2,\perp})\over\sqrt{2}\beta_{1}}-2k_{1}^{+}ie_{\bar{q}}{\psi_{q\bar{q}}(\beta_{1},\beta_{2},\mathbf{k}_{1,\perp},\mathbf{k}_{2,\perp})\over\sqrt{2}\beta_{1}}\right]$
×$\times{\psi_{VR}(x_{V},\mathbf{k}_{R,\perp},x_{R},\mathbf{k}_{V,\perp})\over x_{V}}$ (25)
@@ III.4 Valence Parton Distribution
Inserting Eq.(25) into the expression of hadronic tensor in Eq.(16) and using relations of Eq.(8) and (9) for the valence quark distribution one obtains:
$f_{V}(x_{B},Q^{2})$ $=$ $\int^{Q^{2}}\frac{dx_{1}d^{2}\mathbf{k}_{1,\perp}}{16\pi^{3}x_{1}}\frac{dx_{2}d^{2}\mathbf{k}_{2,\perp}}{16\pi^{3}x_{2}}\frac{dx_{R}d^{2}\mathbf{k}_{R,\perp}}{16\pi^{3}x_{R}}\delta(x_{1}-x_{B})16\pi^{3}\delta\left(1-x_{1}-x_{2}-x_{R}\right)$ (26)
$\times$ $\delta^{(2)}\left(\sum_{i=1,2,R}\mathbf{k}_{i,\perp}\right)\mid\Psi_{q\bar{q}}(\beta_{1},k_{1,\perp},\beta_{2},k_{2,\perp})\mid^{2}\mid\Psi_{VR}(x_{R},k_{R,\perp})\mid^{2}$
Expecting that Light-Front wave function of the $q\bar{q}$ system will depend on relative transverse momentum and relative light-front momentum fractions only, we can factorize above integrals by introducing transfer momentum of quark and anti-quark in the center of mass of the $q\bar{q}$ system
[…]
@@
×$\ \ \ \ \ \ \times\left[\int^{Q^{2}}d^{2}\tilde{\mathbf{k}}_{1,\perp}d^{2}\tilde{\mathbf{k}}_{2,\perp}\delta^{2}(\tilde{\mathbf{k}}_{1,\perp}+\tilde{\mathbf{k}}_{2,\perp})|\Psi_{q\bar{q}}(\beta_{1},\tilde{\mathbf{k}}_{1,\perp})|^{2}\int^{Q^{2}}d^{2}\mathbf{k}_{R,\perp}|\psi_{VR}(x_{R},\mathbf{k}_{R,\perp})|^{2}\right]$
Integrating now by $d^{2}\tilde{k}_{2,\perp}$ using the $\delta^{2}()$ function and renaming $\beta_{1}\equiv\beta$ and $k_{1,\perp}\equiv k_{\perp}$ one obtains:
@@
×$\ \ \ \ \ \ \times\left[\int^{Q^{2}}d^{2}\mathbf{k}_{\perp}|\Psi_{q\bar{q}}(\beta,{\mathbf{k}}_{\perp})|^{2}\right]_{q\bar{q}}\left[\int^{Q^{2}}d^{2}\mathbf{k}_{R,\perp}|\psi_{VR}(x_{R},\mathbf{k}_{R,\perp})|^{2}\right]_{VR}$
where $[\cdots]_{q\bar{q}}$ and $[\cdots]_{VR}$ we can interpret as response functions of valence - cluster and residual systems respectively.
@@ IV.1 General approach for two-body relativistic LFWF
To model $q\bar{q}$ cluster and $V-R$ wave functions we first note that the relativistic normalization of such a wave function is:
$\int\mid\Psi^{LF}(k_{12})\mid^{2}{dy\over y(1-y)}{d^{2}k_{12,\perp}\over 16\pi^{3}}=1$ (30)
@@
$y={k_{1}^{+}\over p_{T}^{+}}$ (31)
with $k_{1}$ and $p_{T}$ being four momenta of one of the bound particles the target particle. Here the relative momentum in the CM of two particles are defined as
@@
$={m_{1}^{2}(1-y)+m_{2}^{2}y+k^{2}_{12,\perp}\over 4y(1-y)}-{m_{1}^{2}+m_{2}^{2}\over 2}+{(m_{1}^{2}-m_{2}^{2})^{2}y(1-y)\over 4(m_{1}^{2}(1-y)+m_{2}^{2}y+k^{2}_{12\perp})}$ (32)
@@
$s_{12}={m_{1}^{2}(1-y)+m_{2}^{2}y+k^{2}_{12,\perp}\over y(1-y)}$ (33)
We model LF wave function similar to Frankfurt and Strikman (1981); Sargsian and Vera (2023) assuming factorization of the relative momentum in the CM system it can be related to the non-relativistic wave function as follows:
$\Psi_{12}^{LF}(k_{12})=\Psi_{12}^{NR}(k_{12})\sqrt{{E_{1}E_{2}\over E_{1}+E_{2}}16\pi^{3}}$ (34)
where non-relativistic wave function is normalized as follows:
$\int\mid\Psi^{NR}(k_{12})\mid^{2}d^{3}k_{12}=1$ (35)

### v2

@@ III.2 Hadronic Current and Hadronic Tensor
$(2\pi)^{4}\delta^{(4)}(p_{\pi}+q-\sum_{i=1}^{2}k_{i}-k_{R})A^{\mu\dagger}A^{\nu},$ (10)
where the sum $\sum_{\{h_{i},\tau_{i}\}}$ is over all the spins of the valence quarks and the residual system. The Lorentz invariant phase space evaluated in light front coordinates is,
@@
$=$ $\frac{dk^{+}d^{2}\mathbf{k}_{\perp}}{2k^{+}}|_{k^{-}=\frac{k_{\perp}^{2}+m^{2}}{k^{+}}}=\frac{dxd^{2}\mathbf{k}_{\perp}}{2x}|_{k^{-}=\frac{k_{\perp}^{2}+m^{2}}{xP_{\pi}^{+}}},$
@@
$\times\delta^{(2)}(\mathbf{p_{\pi}^{\perp}}+\mathbf{q}^{\perp}-\mathbf{k^{\prime\perp}_{1}}-\mathbf{k_{2}^{\perp}}-\mathbf{k}_{R}^{\perp}).$
Taking into account our choice of the reference frame Eq.(7) according to which $q^{+}=0$ and $k^{\prime+}_{1}=k_{1}^{+}$ one obtains:
@@
where $x_{1}={k_{1}^{+}\over p_{\pi}^{+}}$, $x_{2}={k_{2}^{+}\over p_{\pi}^{+}}$, and $x_{R}={k_{R}^{+}\over p_{\pi}^{+}}$. Using $k_{1}^{\prime}=p_{\pi}-k_{2}-k_{3}-k_{R}$ and neglecting initial transverse momenta of pion constituent, compared to $q_{\perp}$,i.e. $k_{1,\perp}^{\prime 2}=(\mathbf{k}_{1,\perp}+\mathbf{q}_{\perp})^{2}\sim\mathbf{q}_{\perp}^{2}\gg k_{1,\perp}^{2}$, as well as using the condition $p_{\pi}^{+}\gg m_{\pi},m_{V},\ m_{R},\ k_{1,2}^{\perp},\ k_{R}$ one evaluates:
@@
$\approx$ $\delta\left(\frac{2p_{\pi}\cdot q}{p_{\pi}^{+}}-\frac{Q^{2}}{x_{1}p_{\pi}^{+}}\right)=\frac{x_{1}p_{\pi}^{+}}{2p_{\pi}\cdot q}\delta(x_{1}-x_{B}),$
where Bjorken $x$ is defined as $x_{B}=\frac{Q^{2}}{2p_{\pi}\cdot q}$. Finally with the same conditions:
$\delta^{(2)}(\mathbf{p_{\pi}^{\perp}}+\mathbf{q}^{\perp}-\mathbf{k^{\prime\perp}_{1}}-\mathbf{k_{2}^{\perp}}-\mathbf{k}_{R}^{\perp})=\delta^{(2)}(\mathbf{k^{\perp}_{1}}+\mathbf{k_{2}^{\perp}}+\mathbf{k}_{R}^{\perp})$ (15)
@@
$W^{\mu\nu}(x,Q^{2})=\frac{1}{4m_{\pi}}\sum_{\{h_{i},\tau_{i}\}}\int[dx][d^{2}\mathbf{k}_{\perp}]{2x_{1}x_{B}\over Q^{2}}\delta(x_{1}-x_{B})A^{\mu\dagger}A^{\nu}$ (16)
@@
$\left[d^{2}\mathbf{k}_{\perp}\right]$ $=$ $16\pi^{3}\delta^{(2)}\left(\sum_{i=1}^{2}\mathbf{k}_{i,\perp}+\mathbf{k}_{R,\perp}\right)\frac{d^{2}\mathbf{k}_{R,\perp}}{16\pi^{3}}\prod_{i=1}^{2}\frac{d^{2}\mathbf{k}_{i,\perp}}{16\pi^{3}}.$ (17)
@@ III.3 Scattering Amplitude in Residual Field Model
To calculate the amplitude presented in Fig.(1)(b) we use light-front diagrammatic rulesLepage and Brodsky (1980). Since momentum transferred to the quark/anti-quark in the pion is significantly larger than their initial momenta, the interference terms will be small and the cross section will represent the incoherent sum of the square of amplitudes corresponding to the scattering from quark and anti-quark. Thus we first calculate amplitude of scattering from quark and then use the result to evaluate the scattering from the anti-quark.
To calculate the amplitude we notice that the diagram of Fig.(1)(b) is characterized by two intermediate states (indicated by vertical dashed lines) that are defined by two light-front propagators. With the momenta assigned for each particles in the scattering defined in the diagram one obtains:
$A^{\mu}(k_{1}^{\prime},h_{1}^{\prime},k_{2},h_{2},k_{R},h_{R})=\sum\limits_{h_{V}}{\bar{u}(k_{1}^{\prime},h_{1}^{\prime})ie_{q}\gamma^{\mu}(/\thinspace k_{1}+m_{q})v(k_{2},h_{2})\Gamma^{V\to q\bar{q}}\chi_{V}\over k_{1}^{+}D_{2}}{\chi_{V}^{\dagger}\chi_{R}^{\dagger}\Gamma^{\pi\to VR}\chi_{\pi}\over k_{V}^{+}D_{1}},$ (18)
where $\Gamma^{\pi\rightarrow VR}$ is the effective vertex for the transition of pion into valence $q\bar{q}$ cluster, $V$ and a residual state $R$ and the $\Gamma^{V\rightarrow q\bar{q}}$ vertex describes the transition of the cluster, $V$ to the valence quark and anti-quark. Also, $h_{i}$ and $h_{1}^{\prime}$ are the helicities of the i’th valence quark before and after the interaction with virtual photon.
For the denominators $D_{1}$ and $D_{2}$ according to light-front diagrammatic rules, one has:
@@
Furthermore, we use LF sum rule for the numerator of the propagator of off-shell quark with momentum $k_{1}$Lepage and Brodsky (1980):
$(/\thinspace k_{1}+m)=\sum\limits_{h_{1}}u(k_{1},h_{1})\bar{u}(k_{1},h_{1})+{\gamma^{+}(k_{1}^{2}-m^{2})\over k_{1}^{+}},$ (21)
and notice that since we are interested in the ++ component of the scattering amplitude, $A^{\mu}$, and due to the relation $\gamma^{+}\gamma^{+}=0$, the second part of Eq.(21) does not contribute to the ++ component of the amplitude. Inserting the first part of Eq.(21) into Eq.(18) one introduces a phenomenological LF wave function for the Valence-Residual system:
@@
and similarly for the $q\bar{q}$ cluster:
$\psi_{q\bar{q}}(\beta_{1},\beta_{2},\mathbf{k}_{1,\perp},\mathbf{k}_{2,\perp})=\frac{\bar{u}(k_{1},h_{1})\Gamma^{V\rightarrow q\bar{q}}v(k_{2},h_{2})\chi_{V}}{m^{2}_{V}-\sum_{i=1}^{2}\frac{k^{2}_{i,\perp}+m^{2}_{i}}{\beta_{i}}},$ (23)
where $\chi_{V}$ and $\chi_{R}$ represent the helicity wave functions for the cluster and residual system respectively.
In further derivation we assume that the residual state has a zero helicity and the quantum number of $q\bar{q}$-cluster is defined by the quantum number of the pion. Writing then the helicity wave function of the cluster explicitly: $\xi_{q\bar{q}}={u(k_{1},h_{1}){\bar{v}}(k_{2},-h_{1})-u(k_{1},-h_{1})\bar{v}(k_{2},h_{1})\over\sqrt{2}}$, for the ”+”-component of the scattering amplitude one obtains:
$A^{+}(k_{1}^{\prime},h_{1}^{\prime},k_{2},-h_{1},k_{R})=\left[\bar{u}(k_{1}^{\prime},h_{1}^{\prime})ie_{q}\gamma^{+}u(k_{1},h_{1}){\psi_{q\bar{q}}(\beta_{1},\beta_{2},\mathbf{k}_{1,\perp},\mathbf{k}_{2,\perp})\over\sqrt{2}\beta_{1}}\right$.
@@
By neglecting the masses of valence quarks, one evaluates only helicity conserving term of the matrix elements of: $\bar{u}(k_{1}^{\prime},h_{1}^{\prime})\gamma^{+}u(k_{1},h_{1})=\bar{v}(k_{1},h_{1})\gamma^{+}v(k_{1}^{\prime},h_{1}^{\prime})=2k_{1}^{+}\delta^{h^{\prime}_{1},h_{1}}$. This results in:
$A^{+}(k_{1}^{\prime},h_{1},k_{2},-h_{1},k_{R})=$ $\left[2k_{1}^{+}ie_{q}{\psi_{q\bar{q}}(\beta_{1},\beta_{2},\mathbf{k}_{1,\perp},\mathbf{k}_{2,\perp})\over\sqrt{2}\beta_{1}}-2k_{1}^{+}ie_{\bar{q}}{\psi_{q\bar{q}}(\beta_{1},\beta_{2},\mathbf{k}_{1,\perp},\mathbf{k}_{2,\perp})\over\sqrt{2}\beta_{1}}\right]$ (25)
$\times{\psi_{VR}(x_{V},\mathbf{k}_{R,\perp},x_{R},\mathbf{k}_{V,\perp})\over x_{V}}.$
@@ III.4 Valence Parton Distribution
Inserting Eq.(25) into the expression of the hadronic tensor in Eq.(16), summing by helicites of final quarks and using relations of Eq.(8) and (9) for the valence quark distribution one obtains:
$f_{V}(x_{B},Q^{2})$ $=$ $\int\frac{dx_{1}d^{2}\mathbf{k}_{1,\perp}}{16\pi^{3}x_{1}}\frac{dx_{2}d^{2}\mathbf{k}_{2,\perp}}{16\pi^{3}x_{2}}\frac{dx_{R}d^{2}\mathbf{k}_{R,\perp}}{16\pi^{3}x_{R}}\delta(x_{1}-x_{B})16\pi^{3}\delta\left(1-x_{1}-x_{2}-x_{R}\right)$ (26)
$\times$ $\delta^{(2)}\left(\sum_{i=1,2,R}\mathbf{k}_{i,\perp}\right)\mid\Psi_{q\bar{q}}(\beta_{1},k_{1,\perp},\beta_{2},k_{2,\perp})\mid^{2}\mid\Psi_{VR}(x_{R},k_{R,\perp})\mid^{2}.$
Expecting that Light-Front wavefunction of the $q\bar{q}$ system will depend on relative transverse momentum and relative light-front momentum fractions only, we can factorize above integrals by introducing transverse momentum of quark and anti-quark in the center of mass of the $q\bar{q}$ system
$\tilde{\mathbf{k}}_{i,\perp}=\mathbf{k}_{i,\perp}-\beta_{i}\mathbf{k}_{V,\perp}\ \ \ \ \text{(i=1,2)}.$ (27)
@@
$\ \ \ \ \ \ \times\left[\int^{Q^{2}}d^{2}\tilde{\mathbf{k}}_{1,\perp}d^{2}\tilde{\mathbf{k}}_{2,\perp}\delta^{2}(\tilde{\mathbf{k}}_{1,\perp}+\tilde{\mathbf{k}}_{2,\perp})|\Psi_{q\bar{q}}(\beta_{1},\tilde{\mathbf{k}}_{1,\perp})|^{2}\int^{Q^{2}}d^{2}\mathbf{k}_{R,\perp}|\psi_{VR}(x_{R},\mathbf{k}_{R,\perp})|^{2}\right].$
The above expression can be integrated by $d^{2}\tilde{k}_{2,\perp}$ using the $\delta^{2}()$ function and renaming $\beta_{1}\equiv\beta$ and $k_{1,\perp}\equiv k_{\perp}$ one obtains:
@@
$\ \ \ \ \ \ \times\left[\int^{Q^{2}}d^{2}\mathbf{k}_{\perp}|\Psi_{q\bar{q}}(\beta,{\mathbf{k}}_{\perp})|^{2}\right]_{q\bar{q}}\left[\int^{Q^{2}}d^{2}\mathbf{k}_{R,\perp}|\psi_{VR}(x_{R},\mathbf{k}_{R,\perp})|^{2}\right]_{VR},$
where $\cdots]_{q\bar{q}}$ and $\cdots]_{VR}$ we can interpret as response functions of the $q\bar{q}$ valence-cluster and residual systems, respectively.
@@ IV.1 General approach for two-body relativistic LFWF
To model $q\bar{q}$ cluster and $VR$ wave functions we first note that both are two-body wave functions for which relativistic normalization is defined as follows:
$\int\mid\Psi^{LF}(k_{12})\mid^{2}{dy\over y(1-y)}{d^{2}k_{12,\perp}\over 16\pi^{3}}=N,$ (30)
@@
$y={k_{1}^{+}\over p_{T}^{+}},$ (31)
with $k_{1}^{\mu}$ and $p_{T}^{\mu}$ being four-momenta of one of the bound particles and the target. Here the relative momentum in the CM frame of two particle system is:
@@
$={m_{1}^{2}(1-y)+m_{2}^{2}y+k^{2}_{12,\perp}\over 4y(1-y)}-{m_{1}^{2}+m_{2}^{2}\over 2}+{(m_{1}^{2}-m_{2}^{2})^{2}y(1-y)\over 4(m_{1}^{2}(1-y)+m_{2}^{2}y+k^{2}_{12\perp})},$ (32)
@@
$s_{12}={m_{1}^{2}(1-y)+m_{2}^{2}y+k^{2}_{12,\perp}\over y(1-y)}.$ (33)
We model LF wave function similar to Frankfurt and Strikman (1981); Sargsian and Vera (2023) assuming factorization of the relative momentum wave function in the CM system and the absence of inelastic transitions in two-body wave function. This allows to relate the CM wave function to the non-relativistic wave function evaluated at the relative momentum defined relativistically in Eq.(32):
$\Psi_{12}^{LF}(k_{12})=\Psi_{12}^{NR}(k_{12})\sqrt{{E_{1}E_{2}\over E_{1}+E_{2}}16\pi^{3}}.$ (34)
Here non-relativistic wave function is normalized as follows:
$\int\mid\Psi^{NR}(k_{12})\mid^{2}d^{3}k_{12}=N.$ (35)
It is worth noting that due to relativistic normalization in Eq.(30) and definition of the light-front relative momentum $k_{12}$ in Eq.(32, the $k_{12}\to\infty$ corresponds to the limits of $y\to 0,1$, thus excluding non-physical region of momenta entering in the argument of LF wave function (as it is the case of non-relativistic wave functions).
The normalization factor $N$, in Eqs.(30) and (35) in general should be fixed through observables, such as baryonic number (or number of constituents)Frankfurt and Strikman (1987); Jaffe and Manohar (1989), or form-factor of the systemSargsian (2010), etc. For example for the case of baryons the valence quark wave function’s normalization can be fixed by the baryonic numberLeon and Sargsian (2022). However, a situation with pion introduces some uncertainty since the baryonic number is zero and this issue will be discussed below.